Ca(OH)2+2HCI=>CaCl2+H2O how many grams of H2O are produced when 86.4 grams of Ca(OH)2 reacts with 89.7 grams of HCI?


1
Expert's answer
2021-10-11T06:12:49-0400

Molar mass of Ca(OH)2= 74.093 g/mol

86.4/74.093= 1.166mol

Molar mass of HCl= 36.458 g/mol

89.7/36.458= 2.46mol

Mole ratio HCl: H2O= 2:1

Mole ratio H2O:Ca(OH)2= 1:1

1.166+1.23=1.396mol

Molar mass of H2O = 18.015

1.396×18.015= 25.15g